An open container holds ice of mass 0.520 kg at a temperature of -13.1 ∘C . The mass of the container can be ignored. Heat is supplied to the container at the constant rate of 900 J/minute . The specific heat of ice to is 2100 J/kg⋅K and the heat of fusion for ice is 334×103J/kg. Part A:How much time tmelts passes before the ice starts to melt? Part B: From the time when the heating begins, how much time trise does it take before the temperature begins to rise above 0∘C?
Answer #1

part A)

let the time taken for the ice to start melting is t1

heat = Power * time = mass of ice * specific heat * change in temperature

900 * t1 = 0.52 * 2100 * (0 - (-13.1))

solving for t1

t1 = 15.9 min

the time taken for the ice to start melting is 15.9 min

part B)

let the time taken for the temperature to rise above 0 degree C is t2

this happens when all the ice melts

heat = Power * time = mass of ice * specific heat * change in temperature + mass * latent heat

900 * t2 = 0.52 * 2100 * (0 - (-13.1)) + 0.52 * 334 *10^3

solving for t2

t2 = 208.9 min

the time taken to increase the temperature from 0 degree C is 208.9 min